the sum of 21 and twice a number is -19. find the number ?
Answer by macston(5194)

Let X be a number
2X+21=-19 Subtract 21 from each side
2x=-40 Divide each side by 2
x=-20: ANSWER The number is -20
CHECK
The sum of 21 and twice -20 is -19
21+2(-20)=-19
21+(-40)=-19
-19=-19